OpenStudy (anonymous):

PLEASSSE HELPP MEEEEEE Match the information about interest earned with the principal amount invested. Remember the equation for calculating simple interest is I equals p times r times t. Match Term Definition $81 over 3 years at an interest rate of 3% $80 over 2 years at an interest rate of 5% $96 dollars over 2 years at an interest rate of 4% $99 dollars over 3 years at an interest rate of 3% A)P=$800 B)P=$1,200 C)P= $900 D)p=$1,100

OpenStudy (anonymous):

PLEASE CAN ANY ONE HELP ME

OpenStudy (dumbcow):

I = Prt solve for P P = I/rt --> 81/.09 = 900 and so on

OpenStudy (mertsj):

What do you need help with? Do you understand the formula you are supposed to use?

OpenStudy (anonymous):

no i dont

OpenStudy (mertsj):

Did you read the problem? The last sentence tells you the formula.

OpenStudy (mertsj):

I=prt

OpenStudy (anonymous):

yes but i suck in math and i just dont get it and this is the last one i have to do im trying to get dome so i can go to sleep

OpenStudy (mertsj):

The fact that you suck in math means you have to work at it harder. In the formula I is the interest. That's what you are supposed to find.

OpenStudy (mertsj):

The way that you find it is by multiplying. What is the principal in the first example in the problem?

OpenStudy (mertsj):

The answer is that you don't know what the principal is. That's what you are supposed to find.

OpenStudy (mertsj):

What is the interest earned?

OpenStudy (anonymous):

900

OpenStudy (mertsj):

$81 over 3 years at an interest rate of 3% That is the first problem and it means that the interest is $81 the time is 3 years and the rate is 3% which is .03 So if you use your formula you have: 81=P(.03)(3)

OpenStudy (mertsj):

Solve that for P

OpenStudy (anonymous):

SO I HAVE TO DO 81X3%

OpenStudy (mertsj):

\[81=P(.03)(3)\] 81=P(.09)

OpenStudy (mertsj):

Do you understand that much?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

\[\frac{81}{.09}=\frac{P(.09)}{.09}\]

OpenStudy (mertsj):

Do you understand that?

OpenStudy (anonymous):

i tried to do the second one but i got 4

OpenStudy (mertsj):

The second one: \[80=P(.05)(2)\]

OpenStudy (anonymous):

is it 800

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

is the 3rd one 1200

OpenStudy (anonymous):

is it ???????

OpenStudy (mertsj):

yes